OpenStudy (babyslapmafro):

Please help. I am asked to find a point P on the line and a vector v parallel to the line by inspection. What is meant by inspection? xi+yj=(2i-j)+t(4i-j)

OpenStudy (babyslapmafro):

Parametric equations: x=2+4t y=-1-t

OpenStudy (babyslapmafro):

P(2,-1) v=<4,-1>